On Monday, 19 October, from 17 - 21 UT, LCO's observatory portal will be closed while we address some software maintenance tasks. You will not be able to submit new observing requests during this time, and the portal website
will show an error message.
The Spectral camera on the 2-meter telescope on Haleakala has been taken out-of-service because of a persistent problem with the camera's shutter. In the past month, the shutter was stuck closed during observations on various nights: Aug 11-15, Aug 30-31, and Sep 6. We have determined the likely source of the malfunction is the shutter control cable. The camera will be returned to service as soon as the shutter problem is fixed.
On August 29, a Sinistro camera, fl12, was installed on the 1-meter telescope in Dome A at Siding Spring (COJ). Although the commissioning observations made since then have been limited (by poor weather), the performance of the fl12 Sinistro exceeds that of fl08, the camera that it replaced. The fl12 Sinistro will be made available for science observations beginning tonight. However, until flat fields have been acquired with the filters required by the science images, only the raw images will be available in the archive. We will begin processing images through the BANZAI pipeline as soon as we have sufficient flat fields.
A version of this article is also available for young scientists! Read it at
The Sinistro camera on the telescope in Dome C at our Sutherland (South Africa) site is now available for science observations. It acquired its first post-commissioning science observations tonight ... before, alas, the dome closed-up because of bad weather. We will continue to monitor this camera's performance, but as always, we welcome your feedback (email firstname.lastname@example.org